# How PARTs assemble into wholes: Learning the relative composition of images

### v1

@@ 5.1 Capabilities of PART
In this section, we start by emphasizing the potential capabilities and opportunities that emerge from adopting relative off-grid sampling in self-supervised learning frameworks, followed by comparisons to other methods. By focusing on these capabilities, we highlight the unique advantages and transformative possibilities that such a paradigm shift offers. Following, we will discuss these four capabilities, namely, (i) off-grid reconstruction, (ii) extension to other aspect ratios and scales, (iii) patch uncertainty, and (iv) symmetry.
Off-grid reconstruction
@@
Unlike grid-based approaches, PART can reconstruct the original image given off-grid patches. This capability is particularly valuable in domains like Satellite and LiDAR imaging, where overlapping patches must be reassembled into the original image. Notably, these patches do not need to originate from a single image. For example, the model can construct a new face by spatially arranging parts of different faces relative to each other. Figure 4 illustrates this capability. The input patches are in the first row and the predictions for different sampling strategies in the second row. In PART, the ground truth input patch visualization consists of a subset of the patches, thus providing a masked input to the model. These images are generated by fixing one random reference patch and positioning all other patches relative to that patch. In the first row, patches are positioned based on the ground truth relative position, and in the second row, patches are positioned based on the model’s output. In the grid sampling strategy, the ground truth relative positions reconstruct the full image because, in this sampling, the patches that form the grid cover the whole image. The model’s prediction almost matches the ground truth, even in small details. The model has learned the general structure of the scenes. For instance, the sky is on top, and the road is at the bottom of the image. It has also learned the triangular structure of the clock. However, some details are missing, such as the hands and the numbers on the clock. It also has difficulties placing mono-color patches because the model only sees the pixel content of the patches. PART’s ability to reconstruct images from off-grid patches demonstrates its understanding of the underlying image structure.
Extension to other aspect ratios and scales
@@
The combination of off-grid patch sampling and relative position prediction as a pretext task, offers the opportunity to reimagine patch sampling in the vision transformers. Once the method can predict relative positions of equally sized patches with an aspect ratio of 1:1, what should stop us from thinking that it could predict the relative positions between patches of varying aspect ratios and scales? In this section, we conduct a proof of concept experiment to show that our method can be extended to include patches of different sizes and aspect ratios. We extend the objective function to include and to capture relative width and height scales:
@@
We also change the sampling to include the coordinates of the bottom right corner of the patches , the patches are then resized to match the patch size of the transformer. We constrain the patch width and height to be at least half the size of the ViT patches and at most double to control having meaningful patches. We pretrain the model using the new objective function for epochs and report its results on the ImageNet classification and COCO Object Detection tasks using the same hyperparameters as our method in Table 1. Extending grid-based methods like MAE to handle multiple aspect ratios and patch sizes is non-trivial. Such an adaptation would require a more sophisticated positional embedding scheme and a decoder architecture capable of upsampling multi-scale latent representations.
@@
Patch uncertainty
Another capability of PART is its ability to estimate patch uncertainties. This is achievable by evaluating whether different reference patches agree in their relative positioning to a single target patch. Our model predicts the relative translation in both scenarios: when patch serves as a reference patch and when it acts as a target patch. We visualize Figure 5 by fixing one target patch and positioning all other patches with respect to that patch. If all patches are placed at the same location, all reference patches agree, thus depicting a more certain patch. Patch uncertainty is particularly valuable in applications such as semantic segmentation for autonomous vehicles or tumor detection in medical imaging, as uncertainty helps determine whether scenarios are common or anomalous.
Symmetry
When a person observes a lip, they naturally anticipate a nose above it. Similarly, upon seeing a nose, they expect a lip below. This demonstrates the concept of symmetry. In our experiment, we aim to demonstrate that PART learns and represents these symmetrical relationships. Figure 6 shows the final prediction matrix of the model versus the ground truth matrix. We can see that the ground truth matrix matches the model prediction for translation in both and axes. The most prominent property that emerges from this figure is the negative symmetry. The negative symmetry is an indication that the model learns that given two patches and with as the reference patch, the model predicts and . Whereas, with as the reference patch, the model predicts and . Considering the heavy masking and lack of global patch information, the model positions two patches correctly relative to each other. Learning the negative symmetry results in consistent relative positioning of object parts, which we expect benefits localization and fine-grained understanding. Especially when samples are scarce, our approach allows for learning from fewer variations due to the built-in symmetries.
5.2 Comparison to grid-based methods
@@
In Table 2, we compare PART with MAE [20], MP3 [21] and DropPos [22] grid-based pretraining methods in the downstream object detection performance. The results demonstrate that PART’s off-grid sampling with overlapping patches and relative patch position prediction improve detection accuracy, particularly for fine-grained local tasks where precise spatial understanding is critical, outperforming methods like MAE, MP3, and DropPos, which rely on fixed grid-based sampling. Notably, while MAE and DropPos use positional embeddings, MP3 and PART do not. PART achieves performance comparable to DropPos, despite DropPos employing additional losses (attentive reconstruction and position smoothing). Without these auxiliary losses, DropPos’s detection performance drops by roughly 2% (Tables 3 & 4 in [22]), highlighting the strength of PART’s objective.
@@
PART can naturally be applied to 1D data. Here we take 1D time series prediction as an example. For 1D data, PART predicts relative time shifts () between randomly-sampled, unequally-spaced windows (the 1D equivalent of off-grid patches) from longer sequences. We validate this approach by pretraining a 1D ViT on biosignals from the PhysioNet 2018 ”You Snooze You Win” Challenge Dataset [76]. As shown in Table 3, our method achieves at least a 2% improvement in sleep staging classification performance compared to both supervised and self-supervised baselines. This task particularly benefits from PART’s capabilities, as accurate sleep staging requires both precise local representations and global understanding of each stage’s position within the complete sequence. Additionally, PART demonstrates superior data efficiency by effectively learning the structure of 1D EEG signals, although limited training data is available.
@@ 5.3 Ablations
| MoCo v3 [65]♯ | ✓ | 300 | 150 | 83.2 |
| DINO [45]♯ | ✓ | 300 | 300 | 82.8 |
| BEiT [56]♯ | ✓ | 800 | 100 | 83.2 |
| CIM [25] | ✓ | 300 | 100 | 83.1 |
@@
Does PART come at the cost of image classification? In Table 6, we compare PART with supervised and state-of-the-art SSL alternatives on the ImageNet-1K [72] classification benchmark. Our method outperforms the supervised results as well as MP3 [21] and shows competitive performance with respect to DropPos [20] and MAE [20]. Note these latter two methods employ position embedding during pretraining. DropPos employs extra position smoothing and attentive reconstruction techniques that could further accelerate training.
@@
We compare PART with supervised and self-supervised alternatives on the CIFAR100 [77] classification benchmark in Table 7. PART consistently outperforms the supervised baselines with and without position embeddings, although it does not use any position embeddings. With only 1000 pretraining epochs, PART outperforms the MP3 [21] baseline with 2000 epochs of pretraining.
@@
As explained in 3.3, a subset is randomly chosen from the patch embeddings. is the parameter that determines the length of . We study the effect of in Figure 8 after 400, 1000, and 4000 epochs of pretraining. We observe that curves follow similar patterns for different epochs of pretraining, while more pretraining epochs result in higher accuracy. We also observe a trade-off in . Higher means the model sees more patch information but must also predict the relative translations for more contradicting patch pairs. Whereas smaller means the model has access to less information, thus overfitting on the task leading to less general representations. There is a sweet spot with patch pairs, where enough global patch information is given to the model, and the training task is neither easy nor difficult.
@@ 6 Discussion & Future Work
So far, we demonstrated the features and capabilities of PART as well as its applicability on multiple data types (1D & 2D), domains (medical & every-day) and tasks (classification & detection). Here, we would like to discuss potential benefits of PART and future directions in more depth.
Hierarchical multi-scale learning: PART’s sampling strategy introduces future research questions such as whether patches should be randomly sampled or should focus on objects or background according to the downstream domain. When extended to multiple scales and aspect ratios, sampled patches of a given object could comprise of parts of that object in different scales leading to hierarchical multi-scale and multi-resolution object representation learning, where objects of different scales must be accurately represented as well as object parts.
Modeling rotations: In the example of when a person observes a lip, they anticipate a nose above it, if the lip is rotated, the expectation is also a rotated nose. How can PART be generalized for rotation equivariance?
@@
Complementary to contrastive learning: PART offers better fine-grained local representations, making it a natural complement to contrastive methods. Together, they can capture both local and global patterns by combining PART’s off-grid relative position prediction with contrastive learning’s view augmentations.
Universal pretraining across diverse data types and domains: PART can be applied to audio spectrograms, video frames and sensor data by adding an additional constraint on time where the model learns a joint relative spatio-temporal relationship. The off-grid sampling allows patches to overlap and vary in size and location, which enables models to learn flexible, continuous representations that better reflect the irregular structures of real-world objects. In domains such as satellite and LiDAR imaging where corresponding images of the same scene overlap, PART can be used for reconstruction. Furthermore, medical imaging with less data and more precise tasks could benefit from this method.
@@
Extension to other tasks: Meanwhile, relative position prediction enhances a model’s ability to reason about spatial relationships between patches in continuous space, improving performance in tasks requiring fine-grained spatial reasoning such as scene graph generation, spatial relation prediction, and 3D reconstruction.
7 Conclusion
The composition of objects and their relative positional relationships provides a rich source of information for representation learning. Building on this insight, we introduced PART, a pretraining method that predicts pairwise relative translations between randomly sampled patches. By employing a random off-grid sampling strategy and relative continuous translation as a pretext task, PART aims to model the relative spatial relationships of objects and the underlying structure of images. Our experiments span 2D and 1D data, where PART’s application indicates a positive impact. Upon finetuning on various downstream tasks such as object detection, time series classification, and image classification, PART has shown promising results compared with existing grid-based supervised and self-supervised baseline methods. Future work could extend the application of PART to more diverse domains and tasks, further refining its capabilities and understanding its full potential.
Acknowledgments
This work was done at Apple and the University of Amsterdam.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used DeepSeek and ChatGPT in order to refine the text. After using these tools, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.

### v2

@@ 5 Results
Having introduced the method, we now analyze the properties learned by PART in practice. We begin with qualitative capabilities that arise from off-grid sampling and relative patch prediction, and then evaluate PART quantitatively against established baselines.
@@ 5.1 Capabilities of PART
In this section, we start by emphasizing the potential capabilities that emerge from adopting relative off-grid sampling in SSL frameworks, followed by comparisons to other methods. These capabilities highlight the unique advantages and transformative possibilities that such a paradigm shift offers, namely, (i) off-grid reconstruction, (ii) extension to other aspect ratios and scales, (iii) patch uncertainty, and (iv) symmetry. For reproducibility, implementation, hyperparameters and the choice of architecture is explained in depth in the Appendix in section 8.1.
(i) Off-grid reconstruction
@@
Unlike grid-based approaches, PART can reconstruct the original image from off-grid patches. This is especially valuable in domains like Satellite and LiDAR imaging, where overlapping patches from different images must be reassembled. These patches need not come from a single image—for instance, the model can compose a new face by arranging parts of different faces. Figure 4 illustrates this: input patches (top row) and the predictions under different sampling strategies (bottom row). In PART, the ground truth visualization consists of a subset of the patches, thus providing a masked input to the model. Images are generated by fixing one random reference patch and positioning all other patches relative to it. In grid sampling, the ground truth positions reconstruct the full image, since patches cover the entire image. The model’s predictions nearly match the ground truth, even in fine details, having learned general scene structure (e.g., sky above, road below, clock’s triangular form). Some details are missing, such as clock hands and numbers, and mono-color patches are harder to place since the model sees only pixel content. PART’s ability to reconstruct from off-grid patches highlights its grasp of underlying image structure.
(ii) Extension to multiple aspect ratios and scales
@@
The combination of off-grid patch sampling and relative position prediction as a pretext task, offers the opportunity to reimagine patch sampling in the vision transformers. If the method can predict relative positions of equal-sized square patches (1:1 aspect ratio), why not also between patches of varying aspect ratios and scales? To test this, we run a proof-of-concept experiment extending the objective to include and for relative width and height:
@@
We modify the sampling to include bottom-right coordinates of the patches, then resizing patches to the ViT patch size . Patch width and height are constrained between half and twice the ViT patch size to ensure meaningful content. The model is pretrained with the new objective for epochs, and results on ImageNet classification and COCO detection are reported in Table 1. Extending grid-based methods like MAE to multiple aspect ratios and patch sizes is non-trivial, requiring more advanced positional embeddings and a decoder that can upsample multi-scale representations.
@@
In practice, the extended model is trained for the same number of epochs and with the same hyperparameters as the base model, without additional tuning. The slight performance drop is likely due to the increased parameter and objective complexity introduced by the additional and terms. We observed in our experiments that the convergence of scale and aspect ratio was fast, while delaying the convergence of and compared to before.
(iii) Patch uncertainty
Another capability of PART is estimating patch uncertainty by checking whether different reference patches agree on the relative position of a target patch. Our model predicts the relative translation both when patch serves as a reference and target patch. In Figure 5, one target patch is positioned relative to all reference patches. If patches are clustered, the model is more certain of that patch. Such uncertainty estimation is valuable in applications like semantic segmentation for autonomous vehicles or tumor detection, where it helps distinguish common from anomalous scenarios.
(iv) Symmetry
When observing a lip, one expects a nose above it; seeing a nose suggests a lip below. This illustrates symmetry. In this experiment, we demonstrate that PART learns and represents these symmetrical relationships. Figure 6 compares the model’s prediction matrix with the ground truth, showing strong alignment along both and axes. The key property that emerges from this figure is negative symmetry: if patch predicts relative to , then predicts relative to . Despite heavy masking and lack of global patch information, the model positions patches correctly relative to each other. Learning the negative symmetry results in consistent relative positioning of object parts, which we expect benefits localization and fine-grained understanding. When samples are scarce, our approach allows for learning from fewer variations due to these built-in symmetries.
5.2 Comparison to grid-based
@@
In Table 2, we compare PART with MAE [mae], MP3 [mp3] and DropPos [droppos] grid-based pretraining methods in the downstream object detection performance. The results demonstrate that PART’s off-grid sampling with overlapping patches and relative patch position prediction improve detection accuracy, particularly for fine-grained local tasks where precise spatial understanding is critical, outperforming methods like MAE, MP3, and DropPos, which rely on fixed grid-based sampling. Notably, while MAE and DropPos use positional embeddings, MP3 and PART do not. PART achieves performance comparable to DropPos, despite DropPos employing additional losses (attentive reconstruction and position smoothing). Without these auxiliary losses, DropPos’s detection performance drops by roughly 2% (Tables 3 & 4 in [droppos]), highlighting the strength of PART’s objective.
@@
PART can also be applied to 1D data. Here we take 1D time series prediction as an example. For 1D data, PART predicts relative time shifts () between randomly-sampled, unequally-spaced windows (the 1D equivalent of off-grid patches) from longer sequences. We validate this approach by pretraining a 1D ViT on biosignals from the PhysioNet 2018 ”You Snooze You Win” Challenge Dataset [physionet2018]. As shown in Table 3, our method achieves at least a 2% improvement in sleep staging classification performance compared to both supervised and self-supervised baselines. This task particularly benefits from PART’s capabilities, as accurate sleep staging requires both precise local representations and global understanding of each stage’s position within the complete sequence. Additionally, as shown in the Appendix 8.2, PART demonstrates superior sample efficiency by effectively learning the structure of 1D EEG signals, although limited training data is available.
@@ 5.3 Ablations
| MoCo v3 [chenempirical]♯ | ✓ | 300 | 150 | 83.2 |
| DINO [dino]♯ | ✓ | 300 | 300 | 82.8 |
| BEiT [beit]♯ | ✓ | 800 | 100 | 83.2 |
| CIM [li2023correlational] | ✓ | 300 | 100 | 83.1 |
@@
Does PART come at the cost of image classification? In Table 6, we compare PART with supervised and state-of-the-art SSL alternatives on the ImageNet-1K [imagenet] classification benchmark. Our method outperforms the supervised results as well as MP3 [mp3] and shows competitive performance with respect to DropPos [mae] and MAE [mae]. Note the latter methods employ position embedding during pretraining. DropPos employs extra position smoothing and attentive reconstruction techniques that could further accelerate training.
@@
We compare PART with supervised and self-supervised alternatives on the CIFAR100 [cifar100] classification benchmark in Table 7. PART consistently outperforms the supervised baselines with and without position embeddings, although it does not use any position embeddings. With only 1000 pretraining epochs, PART outperforms the MP3 [mp3] baseline with 2000 epochs of pretraining.
@@
As explained in 4, a subset is randomly chosen from the patch embeddings. is the parameter that determines the length of . We study the effect of in Figure 8 after 400, 1000, and 4000 epochs of pretraining. We observe that curves follow similar patterns for different epochs of pretraining, while more pretraining epochs result in higher accuracy. We also observe a trade-off in . Higher means the model sees more patch information but must also predict the relative translations for more contradicting patch pairs. Whereas smaller means the model has access to less information, thus overfitting on the task leading to less general representations. There is a sweet spot with patch pairs, where enough global patch information is given to the model, and the training task is neither easy nor difficult.
6 Conclusion
The composition of objects and their parts, along with their relative positions, offers rich information for representation learning. We introduced PART, a pretraining method that predicts continuous relative transformations between random off-grid patches, learning the relative composition of images that generalize beyond occlusions and deformations. We demonstrated PART’s capabilities—off-grid reconstruction, flexible patch forms, patch uncertainty, and symmetry—and how these support the quantitative results. On tasks requiring precise spatial understanding, such as object detection and time-series prediction, PART outperforms grid-based methods like MAE and DropPos, while remaining competitive on global classification. Our experiments show PART’s applicability across data types, domains, and tasks, with potential for further extensions discussed in the next section.
@@ 7 Discussion & Future Work
So far, we demonstrated the capabilities of PART as well as its applicability on multiple data types (1D & 2D), domains (medical & every-day) and tasks (classification & detection). Here, we discuss potential benefits and future directions in depth.
Complementary to contrastive learning:
@@
PART provides fine-grained local representations, making it a complement to contrastive methods. Combined, they can capture both local and global patterns by uniting PART’s off-grid position prediction with contrastive learning’s view augmentations.
Hierarchical multi-scale learning:
PART’s sampling strategy raises questions on whether patches should be sampled randomly or focus on objects or background depending on the downstream task. Extending to multiple scales and aspect ratios could enable hierarchical multi-resolution representations, where objects and their parts at different scales are accurately captured.
Modeling rotations:
For example, seeing a lip suggests a nose above it—but if the lip is rotated, the expectation is a rotated nose. A key question is how PART can be generalized for rotation equivariance.
Extension to other tasks:
@@
Relative position prediction strengthens spatial reasoning in continuous space, benefiting tasks that demand fine-grained spatial understanding such as scene graph generation, spatial relation prediction, and 3D reconstruction.
Universal pretraining across diverse data types and domains:
PART can be extended to audio spectrograms, videos, and sensor data by adding a temporal constraint, learning relative spatio-temporal relationships. Off-grid sampling with overlapping, variable-sized patches enables flexible representations that capture real-world structures. This makes PART useful for reconstruction in satellite and LiDAR imaging, as well as for data-scarce, high-precision domains like medical imaging.
8 Appendix
8.1 Implementation details
For both CIFAR-100 and ImageNet-1k, our pretraining and finetuning configurations closely align with the pipeline outlined in MP3 [mp3], which itself builds upon the foundational work of [vitb]. We adopt the MP3 codebase as our starting point for implementation. During the finetuning phase, we adhere strictly to the supervised training protocols recommended in DeiT [vitb], ensuring consistency with established practices. Detailed descriptions of the implementation specifics for each task are provided below:
Object detection
We evaluate the transfer learning capacity of our method on the COCO Dataset. We perform self-supervised pretraining on the ImageNet-1K [imagenet] with a resolution of 224×224 using ViT-B [vitb] as the backbone. The model is pretrained for epochs with a learning rate of , a batch size of with patch pairs on GPUs. We perform end-to-end finetuning on COCO [COCO2014] for object detection. Specifically, Mask R-CNN [8237584] is finetuned with 1 schedule (12 epochs) and 10241024 resolution. We use the configuration of ViTDet [Li2022ExploringPV] and take ViTB/16 [vit] as the backbone. APb is reported as the main performance metric for object detection. We additionally report AP and AP that correspond to Average Precision at IoU thresholds of 0.5 and 0.75, respectively, following standard COCO evaluation protocols.
1D time series classification
PhysioNet 2018 ”You Snooze You Win” Challenge Dataset [physionet2018] contains multi-channel biosignals that are continuously monitored overnight during sleep studies conducted at Massachusetts General Hospital. The task is to predict one of five sleep stages (Wake, Non-REM1, Non-REM2, Non-REM3, REM) given a 30-second window of data containing six channels of scalp electroencephalography (EEG). The EEG data is bandpass filtered with cutoffs 0.1-30 Hz and then re-sampled to have a 100 Hz sampling rate. Thirty-second windows are instance normalized as an additional pre-processing step before being tokenized by a linear layer. Forty 1-second patches are then randomly sampled. We use recordings from 1,653 subjects across the entire dataset for all pretraining strategies. For finetuning, we down-sample the dataset to 10 subjects to simulate a low-labeled data regime. For testing, the finetuned models are evaluated on recordings from 200 subjects, which are held out from both pretraining and finetuning stages. We randomly mask out 20% of the position embeddings for this experiment. All experiments use a 1D ViT backbone with 12.9M parameters and the input to the model is the 1D data for both pretraining and finetuning.
Image classification
We perform self-supervised pretraining on CIFAR100 [cifar100] and ImageNet-1K [imagenet] with a resolution of 32x32 and 224×224 respectively. Following Zhai et al. [mp3], we use ViT-S as the backbone of CIFAR100 and ViT-B [vitb] as the backbone of ImageNet-1K. During pretraining, we perform a hyperparameter search on learning rates , and the number of pairs and choose the best result for each experiment. We perform epochs of finetuning on CIFAR100 and epochs of finetuning on ImageNet-1K with a learning rate of . We report accuracy, error, and the mean squared error in and dimensions.
Choice of Architecture
8.2 Sample efficiency
To evaluate sample efficiency, we varied the number of subjects used for fine-tuning in the EEG sleep staging task from 10 to 657, and repeated 5 times with different random seeds and sample subsets. As shown below, PART consistently outperforms baselines across all label fractions with fewer fine-tuning samples.
| # Fine-tuning Samples | MP3 | DropPos | MAE | PART |
| 10 | 0.56 | 0.58 | 0.62 | 0.64 |
| 50 | 0.63 | 0.64 | 0.65 | 0.67 |
| 100 | 0.65 | 0.66 | 0.67 | 0.68 |
| 657 | 0.68 | 0.68 | 0.68 | 0.70 |